Lesson 14 question number 4.Solutions of iron (III) chloride, FeCl3, are used in photoengraving and to make ink. This compound can be made by the following reaction: 2Fe + 3Cl2 = 2FeCl3

a. How many grams of Fe are needed to combine with 4.5 moles of Cl2?

b. If 240 g of Fe is to be used in this reaction, with adequate Cl2, how many moles of FeCl3 will be produced?

5.Ammonia is produced synthetically by the reaction below. How many moles of NH3 are formed when 200.0 g of N2reacts with hydrogen?

N2 + 3H2 = 2NH3

6.Iron metal is produced in a blast furnace by the reaction of iron (III) oxide and coke (pure carbon). If 25.0 moles of pure Fe2O3 is used, how many grams of iron can be produced? The balanced chemical equation for the reaction is: Fe2O3 + 3C = 2Fe + 3CO



1
Expert's answer
2021-03-22T06:29:18-0400

Solution:

(Q4a):

The balanced chemical equation:

2Fe + 3Cl2 → 2FeCl3

According to the equation: n(Fe)/2 = n(Cl2)/3

n(Fe) = 2 × n(Cl2) / 3 = (2 × 4.5 mol) / 3 = 3.0 mol


Mass of Fe = Moles of Fe × Molar mass of Fe

The molar mass of Fe is 55.845 g mol-1.

Hence,

Mass of Fe = 3.0 mol × 55.845 g mol-1 = 167.535 g = 167.5 g

Mass of Fe = 167.5 g


OR (short form solution):

(4.5 mol Cl2) × (2 mol Fe / 3 mol Cl2) × (55.845 g Fe / 1 mol Fe) = 167.5 g Fe


(Q4b):

The balanced chemical equation:

2Fe + 3Cl2 → 2FeCl3

According to the equation: n(Fe) = n(FeCl3)


Moles of Fe = Mass of Fe / Molar mass of Fe

The molar mass of Fe is 55.845 g mol-1.

Hence,

Moles of Fe = 240 g / 55.845 g mol-1 = 4.2976 mol


n(FeCl3) = n(Fe) = 4.2976 mol = 4.3 mol

Moles of FeCl3 = 4.3 mol


OR (short form solution):

(240 g Fe) × (1 mol Fe / 55.845 g Fe) × (2 mol FeCl3 / 2 mol Fe) = 4.3 mol FeCl3


(Q5):

The balanced chemical equation:

N2 + 3H2 → 2NH3

According to the equation: n(N2) = n(NH3)/2


Moles of N2 = Mass of N2 / Molar mass of N2

The molar mass of N2 is 28.013 g mol-1.

Hence,

Moles of N2 = 200.0 g / 28.013 g mol-1 = 7.1395 mol


n(NH3) = 2 × n(N2) = 2 × 7.1395 mol = 14.279 mol = 14.28 mol

Moles of NH3 = 14.28 mol


OR (short form solution):

(200.0 g N2) × (1 mol N2 / 28.013 g N2) × (2 mol NH3 / 1 mol N2) = 14.28 mol NH3


(Q6):

The balanced chemical equation:

Fe2O3 + 3C → 2Fe + 3CO

According to the equation: n(Fe2O3) = n(Fe)/2

n(Fe) = 2 × n(Fe2O3) = 2 × 25.0 mol = 50.0 mol


Mass of Fe = Moles of Fe × Molar mass of Fe

The molar mass of Fe is 55.845 g mol-1.

Hence,

Mass of Fe = 50.0 mol × 55.845 g mol-1 = 2792.25 g

Mass of Fe = 2792.25 g


OR (short form solution):

(25.0 mol Fe2O3) × (2 mol Fe / 1 mol Fe2O3) × (55.845 g Fe / 1 mol Fe) = 2792.25 g Fe


Answers:

(Q4a): 167.5 g Fe

(Q4b): 4.3 mol FeCl3

(Q5): 14.28 mol NH3

(Q6): 2792.25 g Fe